### v1

@@ 0.2.3 Previous NLP Verification Approaches
The main advantage of randomised smoothing-based methods is their scalability, indeed recent approaches are tested on larger transformer such as BERT and Alpaca. However, their main issue is that they are probabilistic approaches, meaning they give certifications up to a certain probability (e.g., 99.9%). In this work we focus on deterministic approaches, hence we only report these works in Table 1 for completeness without delving deeper into each paper here. All randomised smoothing-based approaches use data augmentation obtained by semantic perturbations.
0.2.4 Data Sets and Use Cases used in NLP Verification
Existing NLP verification data sets.
Table 2 summarises the main features and tasks of the datasets used in NLP verification. Despite their diverse origins and applications, the use of these datasets in the NLP verification literature converge on text classification, with a predominant focus on binary or multi-class categorisation. Furthermore, datasets can be sensitive to perturbations, i.e. perturbations can have non-trivial impact on label consistency. For example, Jia et al. [54] use IBP with the SNLI [15] [^3] A semantic inference dataset that labels whether one sentence entails, contradicts or is neutral to another sentence. dataset (see Tables 1 and 2) to show that word perturbations (e.g. ‘good’ to ‘best’) can change whether one sentence entails another. Some works such as [54] try to address this label consistency, while others do not.
Additionally, we find that the previous research on NLP verification does not utilise safety critical datasets (which strongly motivates the choice of datasets in alternative verification domains), with the exception of [37]. The papers do not provide detailed motivation as to why the dataset choices were made, however it could be due to the datasets being commonly used in NLP benchmarks (IMDB …). Instead, we focus on datasets motivated by safety critical applications.
Dataset Safety Critical Category Tasks Size Classes
IMDB [84]
$\times$
Sentiment analysis
Document-level and sentence-level classification
25,000
2
SST [112]
$\times$
Sentiment analysis
Sentiment classification, hierarchical sentiment classification, sentiment span detection
70,042
5
SST2 [112]
$\times$
Sentiment analysis
Sentiment classification
70,042
2
YELP [106]
$\times$
Sentiment analysis
Sentiment classification
570,771
2
Rotten Tomatoes Movie Review [96]
$\times$
Sentiment analysis
Sentiment classification
48,869
3/4
Amazon [86]
$\times$
Sentiment analysis
Sentiment classification, aspect-based sentiment analysis
34,686,770
5
SNLI [15]
$\times$
Semantic inference
Natural language inference, semantic similarity
570,152
3
MNLI [127]
$\times$
Semantic inference
Natural language inference, semantic similarity, generalisation
432,702
3
AGNews [143]
$\times$
Text analysis
Text classification, sentiment classification
127,600
4
CogComp QC [74]
$\times$
Text analysis
Question classification, semantic understanding
15,000
6/50
Toxic Comment [26]
Text analysis
Toxic comment classification, fine-grained toxicity analysis, bias analysis
18,560
6
Data Sets Proposed in This Paper
In this paper we propose to use two datasets that have not been used in NLP verification literature before. Both are driven by real-world use cases of safety-critical NLP applications, i.e. applications for which law enforcement and safety demand formal guarantees of “good" DNN behaviour.
Chatbot Disclosure Dataset.
First case concerns new legislation which states that a chatbot must not mislead people about its artificial identity [68, 63]. Given that the regulatory landscape surrounding NLP models (particularly LLMs and generative AI) is rapidly evolving, similar legislation could be widespread in the future – with recent calls for the US Congress to formalise such disclosure requirements [90]. The prohibition on deceptive conduct act may apply to the outputs generated by NLP systems if used commercially [4], and at minimum a system must guarantee a truthful response when asked about its agency [47, 1]. Furthermore, the burden of this should be placed on the designers of NLP systems, and not on the consumers.
Our first safety critical case is the R-U-A-Robot dataset [47], a written English dataset consisting of 6800 variations on queries relating to the intent of ‘Are you a robot?’, such as ‘I’m a man, what about you?’. The dataset was created via a context-free grammar template, crowd-sourcing and pre-existing data sources. It consists of 2,720 positive examples (where given the query, it is appropriate for the system to state its non-human identity), 3,400 negative/adversarial examples and 680 ‘ambiguous-if-clarify’ examples (where it is unclear whether the system is required to state its identity). The dataset was created to promote transparency which may be required when the user receives unsolicited phone calls from artificial systems. Given systems like Google Duplex [70], and the criticism it received for human-sounding outputs [77], it is also highly plausible for the user to be deceived regarding the outputs generated by other NLP-based systems [4]. Thus we choose this dataset to understand how to enforce such disclosure requirements. We collapse the positive and ambiguous examples into one label, following the principle of ‘better be safe than sorry’, i.e. prioritising a high recall system.
Medical Safety Dataset.
Another scenario one might consider is that inappropriate outputs of NLP systems have the potential to cause harm to human users [12]. For example, a system may give a user false impressions of its ‘expertise’ and generate harmful advice in response to medically related user queries [34]. In practice it may be desirable for the system to avoid answering such queries. Thus we choose the Medical safety dataset [2], a written English dataset consisting of 2,917 risk-graded medical and non medical queries (1,417 and 1,500 examples respectively). The dataset was constructed via collecting questions posted on reddit, such as r/AskDocs. The medical queries have been labelled by experts and crowd annotators for both relevance and levels of risk (i.e. non-serious, serious to critical) following established World Economic Forum (WEF) risk levels designated for chatbots in healthcare [130]. We merge the medical queries of different risk-levels into one class, given the high scarcity of the latter 2 labels to create an in-domain/out-of-domain classification task for medical queries. Additionally, we consider only the medical queries that were labelled as such by expert medical practitioners. Thus this dataset will facilitate discussion on how to guarantee a system recognises medical queries, in order to avoid generating medical output.
0.2.5 Our Work (NLP Verification Pipelines)
To show relation of our work to the body of already existing work, we distill an “NLP verification pipeline” that is common across many related papers. Figure 2 shows the pipeline diagrammatically. It proceeds in stages:
@@ 0.2.5 Our Work (NLP Verification Pipelines)
    Given an NLP dataset, generate semantic perturbations on sentences that it contains. The semantic perturbations can be of different kinds: character, word or sentence level. IBP and randomised smoothing use word and character perturbations, abstract interpretation papers usually do not use any semantic perturbations. Our method allows to use all existing semantic perturbations, in particular, we implement character and word level perturbations as in [91], sentence level perturbations with PolyJuice [132] and Vicuna.
@@
    Given a notion of geometric/semantic subspaces, one can use it in two different ways:
    • 4.
@@
    • 5.
      $\ldots$ verify the classifier’s behaviour within those subspaces. IBP papers use IBP algorithms which are incomplete and imprecise, abstract interpretation gives incomplete and imprecise methods, and we use SMT-based tool Marabou (complete and precise) and abstract-interpretation tool ERAN (incomplete and imprecise).
Table 1 summarises differences and similarities of the above NLP verification approaches against ours. To the best of our knowledge, we are the first to use complete and precise methods in NLP verification. This paper is the first to employ SMT-based verifiers and to show how they out-perform abstract interpretation-based verification, which produce tighter bounds than IBP-based methods.
Furthermore, our study is the first to demonstrate that the construction of semantic subspaces can happen independently of the choice of the training and verification algorithms. Likewise, although training and verification build upon the defined (semantic) subpaces, the actual choice of the training and verification algorithms can be made independently of the method used to define the semantic subspaces. This separation, and the general modularity of our approach, facilitates a comprehensive examination and comparison of the two key components involved in any NLP verification process:
@@
  • relation between the volume/shape of semantic subpaces and verifiability of neural networks obtained via semantic training with these subpaces.
@@
0.3 The Parametric NLP Verification Pipeline
@@
This section presents a parametric NLP verification pipeline, shown in Figure 2 diagrammatically. We call it “parametric” because each component within the pipeline operates independently of the others and can be taken as a parameter when studying other components. The parametric nature of the pipeline allows for the seamless integration of state-of-the-art methods at every stage, and for more sophisticated experiments with those methods. Note that the outlined pipeline can be seen as a filter which can be applied on top of an NLP system or LLM (such as S-BERT and S-GPT) to certify intended DNN behavior for safety-critical input queries. The following section provides a detailed exposition of the methodological choices made at each step of the pipeline.
0.3.1 Semantic Perturbations
As discussed in Section 0.2.5, we require semantic perturbations for creating semantic subspaces. To do so, we consider three kinds of perturbations – i.e. character, word and sentence level. This systematically accounts for different variations of the samples.
Character and word level perturbations
are created via a rule-based method proposed in [91], to simulate different kinds of noise one could expect from spelling mistakes, typos etc. These perturbations are non-adversarial and can be generated automatically. The authors found that NLP models are sensitive to such small errors, while in practice this should not be the case. Character level perturbations types include randomly inserting, deleting, replacing, swapping or repeating a character of the data sample. At the character level, we do not apply letter case changing, given it does not change the sentence-level representation of the sample, nor do we apply perturbations to commonly misspelled words, given only a small overlap of those words occur in our datasets. Perturbations types at the word level include randomly repeating or deleting a word, changing the ordering of the words, the verb tense, singular verbs to plural verbs or adding negation to the data sample. At the word level, we omit replacement with synonyms, as this is accounted for via sentence rephrasing. Negation is not done on the medical safety dataset, as it creates label ambiguities (e.g. ‘pain when straightening knee’ $\rightarrow$ ‘no pain when straightening knee’), as well as singular plural tense and verb tense, given human annotators would experience difficulties with this task (e.g. rephrase the following in plural/ with changed tense – ‘peritonsillar abscess drainage aftercare.. please help’).
Further examples of character and word rule-based perturbations can be found in Tables 3 and 4.
Method Description Original sentence Altered sentence
Insertion
A character is randomly selected and inserted in a random position.
Are you a robot?
Are yovu a robot?
Deletion
A character is randomly selected and deleted.
Are you a robot?
Are you a robt?
Replacement
A character is randomly selected and replaced by an adjacent character on the keyboard.
Are you a robot?
Are you a ronot?
Swapping
A character is randomly selected and swapped with the adjacent right or left character in the word.
Are you a robot?
Are you a rboot?
Repetition
A character in a random position is selected and duplicated.
Are you a robot?
Arre you a robot?
@@ 0.3.1 Semantic Perturbations
Method Description Original sentence Altered sentence
Deletion
Randomly selects a word and removes it.
Can u tell me if you are a chatbot?
Can u tell if you are a chatbot?
Repetition
Randomly selects a word and duplicates it.
Can u tell me if you are a chatbot?
Can can u tell me if you are a chatbot?
Negation
Identifies verbs then flips them (negative/positive).
Can u tell me if you are a chatbot?
Can u tell me if you are not a chatbot?
Singular/ plural verbs
Changes verbs to singular form, and conversely.
Can u tell me if you are a chatbot?
Can u tell me if you is a chatbot?
Word order
Randomly selects consecutive words and changes the order in which they appear.
Can u tell me if you are a chatbot?
Can u tell me if you are chatbot a?
Verb tense
Converts present simple or continuous verbs to their corresponding past simple or continuous form.
Can u tell me if you are a chatbot?
Can u tell me if you were a chatbot?
@@
Sentence level perturbations.
We experiment with two types of sentence level perturbations, particularly due to the complicated nature of the medical queries (e.g. it is non-trivial to rephrase queries such as this – ‘peritonsillar abscess drainage aftercare.. please help’). We do so by either using Polyjuice [132] or vicuna-13b [^4] Using the following API: https://replicate.com/replicate/vicuna-13b/api.. Polyjuice is a general-purpose counterfactual generator that allows for control over perturbation types and locations, trained by fine-tuning GPT-2 on multiple datasets of paired sentences. Vicuna is a state-of-the-art open source chatbot trained by fine-tuning LLaMA [115] on user-shared conversations collected from ShareGPT  [^5] https://sharegpt.com/. For Vicuna, we use the following prompt to generate variations on our data samples ‘Rephrase this sentence 5 times: “[Example]”.’ For example, from the sentence “How long will I be contagious?", we can obtain “How many years will I be contagious?" or “Will I be contagious for long?" and so on.
@@
0.3.2 NLP Embeddings
@@
Next component of the pipeline are embeddings. Embeddings play a crucial role in NLP verification as they map textual data into continuous vector spaces, in a way that should capture semantic relationships and contextual information.
@@
Given an NLP dataset $\mathcal{Y}$ as a set of sentences $s_{1},\ldots,s_{q}$ written in natural language, an embedding $\mathbb{E}$ is a function that maps a sentence to a vector in ${}^{m}$. The vector space ${}^{m}$ is called the embedding space. Ideally, $\mathbb{E}$ should reflect the semantic similarities between sentences in $\mathcal{Y}$, i.e. the more semantically similar two sentences $s_{i}$ and $s_{j}$ are, the closer the distance between $\mathbb{E}(s_{i})$ and $\mathbb{E}(s_{j})$ should be in ${}^{m}$. Of course, defining semantic similarity in precise terms may not be tractable (the number of unseen sentences maybe infinite, the similarity maybe understood subjectively and/or depending on the context). This is why, the state-of-the-art NLP relies on machine learning methods to capture the notion of semantic similarity approximately.
@@
Currently, the most common approach to obtain an embedding function $\mathbb{E}$ is by training transformers [32, 101]. Transformers are a type of DNNs that can be trained to map sequential data into real vector spaces and are capable of handling variable-length input sequences. They can also be used for other tasks, such as classification or sentence generation, but in those cases, too, training happens at the level of embedding spaces. In this work, a transformer is trained as a function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$ for some given $m$. The key feature of the transformer is the “self-attention mechanism", which allows the network to weigh the importance of different elements in the input sequence when making predictions, rather than relying solely on the order of elements in the sequence. This makes them good at learning to associate semantically similar words or sentences. In this work we initially use Sentence-BERT [101] and later add Sentence-GPT [92] to embed sentences. Unfortunately, the relation between the embedding space and the NLP dataset is not bijective: i.e. each sentence is mapped into the embedding space, but not every point in the embedding space has a corresponding sentence. This problem is well-known in NLP literature [65] and, as shown in this paper, is one of the reasons why verification of NLP is tricky. Given an NLP dataset $\mathcal{Y}$ that should be classified into $n$ classes, the standard approach is to construct a function $N:^{m}\rightarrow^{n}$ that maps the embedded inputs to the classes. In order to do that, a domain specific classifier $N$ is trained on the embeddings $\mathbb{E}(\mathcal{Y})$ and the final system will then be the composition of the two subsystems, i.e. $N\circ\mathbb{E}$.
0.3.3 Geometric Analysis of Embedding Spaces
We now formally define geometric and semantic subpaces of the embedding space. Our goal is to define subpaces on the embedding space ${}^{m}$ by using an effective algorithmic procedure.

### v3

@@ 2.6 Previous NLP Verification Approaches
The main advantage of randomised smoothing-based methods is their scalability, indeed recent approaches are tested on larger transformer such as BERT and Alpaca. However, their main issue is that they are probabilistic approaches, meaning they give certifications up to a certain probability. In this work we focus on deterministic approaches, hence we only report these works in Table 2 for completeness without delving deeper into each paper here. All randomised smoothing-based approaches use data augmentation obtained by semantic perturbations.
To systematically compare the existing body of research, we distil an “NLP verification pipeline” that is common across many related papers. This pipeline is outlined diagrammatically in Figure 2, while Tables 1 and 2 provide a detailed breakdown, with columns corresponding to each stage of the pipeline. It proceeds in stages:
@@
    Given an NLP dataset, generate semantic perturbations on sentences that it contains. The semantic perturbations can be of different kinds: character, word or sentence level. IBP and randomised smoothing use word and character perturbations, abstract interpretation papers usually do not use any semantic perturbations. Tables 1 and 2 give the exact mapping of perturbation methods to papers. Our method allows to use all existing semantic perturbations, in particular, we implement character and word level perturbations as in Moradi et al. [103], sentence level perturbations with PolyJuice [21] and Vicuna.
@@
  4. 4.
@@
  5. 5.
    Use the geometric/semantic subspaces to verify the classifier’s behaviour within those subspaces. The papers [17, 18, 107, 19, 108] use IBP algorithms and the papers  [109, 15, 110, 111] use abstract interpretation; in both cases it is incomplete and deterministic verification. See ‘Verification algorithm’ and ‘Verification characteristics’ columns of Tables 1 and 2. We use SMT-based tool Marabou (complete and deterministic) and abstract-interpretation tool ERAN (incomplete and deterministic).
Tables 1 and 2 summarise differences and similarities of the above NLP verification approaches against ours. To the best of our knowledge, we are the first to use SMT-based complete methods in NLP verification and we show how they achieve higher verifiability than abstract interpretation verification approaches (ERAN and CORA) or IBP and BaB ($\alpha\beta$-CROWN), thanks to the increased precision of the ReLUplex algorithm that underlies Marabou.
Furthermore, our study is the first to demonstrate that the construction of semantic subspaces can happen independently of the choice of the training and verification algorithms. Likewise, although training and verification build upon the defined (semantic) subspaces, the actual choice of the training and verification algorithms can be made independently of the method used to define the semantic subspaces. This separation, and the general modularity of our approach, facilitates a comprehensive examination and comparison of the two key components involved in any NLP verification process:
@@
  • relation between the volume/shape of semantic subspaces and verifiability of neural networks obtained via semantic training with these subspaces.
@@
3 The Parametric NLP Verification Pipeline
This section presents a parametric NLP verification pipeline, shown in Figure 2 diagrammatically. We call it “parametric” because each component within the pipeline is defined independently of the others and can be taken as a parameter when studying other components. The parametric nature of the pipeline allows for the seamless integration of state-of-the-art methods at every stage, and for more sophisticated experiments with those methods. The following section provides a detailed exposition of the methodological choices made at each step of the pipeline.
@@ 3 The Parametric NLP Verification Pipeline
3.1 Semantic Perturbations
As discussed in Section 2.6, we require semantic perturbations for creating semantic subspaces. To do so, we consider three kinds of perturbations – i.e. character, word and sentence level. This systematically accounts for different variations of the samples.
Character and word level perturbations are created via a rule-based method proposed by Moradi et al. [103] to simulate different kinds of noise one could expect from spelling mistakes, typos etc. These perturbations are non-adversarial and can be generated automatically. Moradi et al. [103] found that NLP models are sensitive to such small errors, while in practice this should not be the case. Character level perturbations types include randomly inserting, deleting, replacing, swapping or repeating a character of the data sample. At the character level, we do not apply letter case changing, given it does not change the sentence-level representation of the sample. Nor do we apply perturbations to commonly misspelled words, given only a small percentage of the most commonly misspelled words occur in our datasets. Perturbations types at the word level include randomly repeating or deleting a word, changing the ordering of the words, the verb tense, singular verbs to plural verbs or adding negation to the data sample. At the word level, we omit replacement with synonyms, as this is accounted for via sentence rephrasing. Negation is not done on the medical safety dataset, as it creates label ambiguities (e.g. ‘pain when straightening knee’ $\rightarrow$ ‘no pain when straightening knee’), as well as singular plural tense and verb tense, given human annotators would experience difficulties with this task (e.g. rephrase the following in plural/ with changed tense – ‘peritonsillar abscess drainage aftercare.. please help’). Note that the Medical dataset contains several sentences without a verb (like the one above) for which it is impossible to pluralise or change the tense of the verb.
Further examples of character and word rule-based perturbations can be found in Tables 4 and 5.
Method Description Altered sentence
(Are you a robot?)
Insertion A character is randomly selected and inserted in a random position. Are yovu a robot?
Deletion A character is randomly selected and deleted. Are you a robt?
Replacement A character is randomly selected and replaced by an adjacent character on the keyboard. Are you a ronot?
Swapping A character is randomly selected and swapped with the adjacent right or left character in the word. Are you a rboot?
Repetition A character in a random position is selected and duplicated. Arre you a robot?
@@ 3.1 Semantic Perturbations
Method Description Altered sentence
(Can u tell me if you are a chatbot?)
Deletion Randomly selects a word and removes it. Can u tell if you are a chatbot?
Repetition Randomly selects a word and duplicates it. Can can u tell me if you are a chatbot?
Negation Identifies verbs then flips them (negative/positive). Can u tell me if you are not a chatbot?
Singular/ plural verbs Changes verbs to singular form, and conversely. Can u tell me if you is a chatbot?
Word order Randomly selects consecutive words and changes the order in which they appear. Can u tell me if you are chatbot a?
Verb tense Converts present simple or continuous verbs to their corresponding past simple or continuous form. Can u tell me if you were a chatbot?
@@
Sentence level perturbations. We experiment with two types of sentence level perturbations, particularly due to the complicated nature of the medical queries (e.g. it is non-trivial to rephrase queries such as this – ‘peritonsillar abscess drainage aftercare.. please help’). We do so by either using Polyjuice [21] or vicuna-13b [^4] Using the following API: https://replicate.com/replicate/vicuna-13b/api.. Polyjuice is a general-purpose counterfactual generator that allows for control over perturbation types and locations, trained by fine-tuning GPT-2 on multiple datasets of paired sentences. Vicuna is a state-of-the-art open source chatbot trained by fine-tuning LLaMA [138] on user-shared conversations collected from ShareGPT  [^5] https://sharegpt.com/. For Vicuna, we use the following prompt to generate variations on our data samples ‘Rephrase this sentence 5 times: “[Example]”.’ For example, from the sentence “How long will I be contagious?”, we can obtain “How many years will I be contagious?” or “Will I be contagious for long?” and so on.
@@
Semantic similarity of perturbations. In later sections we will make an assumption that the perturbations that we use produce sentences that are semantically similar to the originals. However, precisely defining or measuring semantic similarity is a challenge in its own right, as semantic meaning of sentences can be subjective, context-dependent, which makes evaluating their similarity intractable. Nevertheless, Subsection 5.5.2 will discuss and use several metrics for calculating semantic similarity of sentences, modulo some simplifying assumptions.
3.2 NLP Embeddings
@@
The next component of the pipeline is the embeddings. Embeddings play a crucial role in NLP verification as they map textual data into continuous vector spaces, in a way that should capture semantic relationships and contextual information.
@@
Given the set of all strings, $\mathbb{S}$, an NLP dataset $\mathcal{Y}\subset\mathbb{S}$ is a set of sentences $s_{1},\ldots,s_{q}$ written in natural language. The embedding $E$ is a function that maps a string in $\mathbb{S}$ to a vector in m. The vector space m is called the embedding space. Ideally, $E$ should reflect the semantic similarities between sentences in $\mathbb{S}$, i.e. the more semantically similar two sentences $s_{i}$ and $s_{j}$ are, the closer the distance between $E(s_{i})$ and $E(s_{j})$ should be in m. Of course, defining semantic similarity in precise terms may not be tractable (the number of unseen sentences may be infinite, the similarity may be subjective and/or depend on the context). This is why, the state-of-the-art NLP relies on machine learning methods to capture the notion of semantic similarity approximately.
@@
Currently, the most common approach to obtain an embedding function $E$ is by training transformers [139, 140]. Transformers are a type of DNNs that can be trained to map sequential data into real vector spaces and are capable of handling variable-length input sequences. They can also be used for other tasks, such as classification or sentence generation, but in those cases, too, training happens at the level of embedding spaces. In this work, a transformer is trained as a function $E:\mathbb{S}\rightarrow^{m}$ for some given $m$. The key feature of the transformer is the “self-attention mechanism”, which allows the network to weigh the importance of different elements in the input sequence when making predictions, rather than relying solely on the order of elements in the sequence. This makes them good at learning to associate semantically similar words or sentences. In this work we initially use Sentence-BERT [140] and later add Sentence-GPT [141] to embed sentences. Unfortunately, the relation between the embedding space and the NLP dataset is not bijective: i.e. each sentence is mapped into the embedding space, but not every point in the embedding space has a corresponding sentence. This problem is well-known in NLP literature [142] and, as shown in this paper, is one of the reasons why verification of NLP is tricky. Given an NLP dataset $\mathcal{Y}$ that should be classified into $n$ classes, the standard approach is to construct a function $N:^{m}\rightarrow^{n}$ that maps the embedded inputs to the classes. In order to do that, a domain specific classifier $N$ is trained on the embeddings $E(\mathcal{Y})$ and the final system will then be the composition of the two subsystems, i.e. $N\circ E$.
3.3 Geometric Analysis of Embedding Spaces
In the recent years, the study of manifold subspaces has gained significant attention in the context of machine learning verification [55], where the geometry of data regions plays an important role. In this section, we formally define most common subspaces used in verification: convex sets, convex hulls, zonotopes, and hyper-rectangles (also known as multi-dimensional intervals), following closely [55].
Definition 1 (Convex Set)
A set $\mathbb{Z}\subseteq^{m}$ is said to be convex if, for any two points $x_{1},x_{2}\in\mathbb{Z}$, the line segment joining them is entirely contained within $\mathbb{Z}$. Formally, this means that for all $x_{1},x_{2}\in\mathbb{Z}$ and $\lambda\in[0,1]$, the points
$\lambda x_{1}+(1-\lambda)x_{2}\in\mathbb{Z}.$
In other words, a set is convex if, for any pair of points in the set, the entire segment connecting them lies within the set.
The convex hull of a set is the smallest convex set that contains all the points of the set. It can be seen as the “tightest” boundary enclosing the points.
Definition 2 (Convex Hull)
The convex hull of a set $\mathbb{Z}\subseteq^{m}$, denoted by $\text{conv}(\mathbb{Z})$, is the smallest convex set containing $\mathbb{Z}$. Formally, it can be defined as:
$\text{conv}(\mathbb{Z}):=\left\{\sum_{i=1}^{p}\lambda_{i}x_{i}\;\middle|\;x_{i}\in\mathbb{Z},\;\lambda_{i}\geq 0,\;\sum_{i=1}^{p}\lambda_{i}=1,\;p\in{\rm Nature}\right\}.$
In other words, $\text{conv}(\mathbb{Z})$ consists of all finite convex combinations of points in $\mathbb{Z}$. The construction of the convex hull has a complexity of $O(p^{m/2})$, where $p$ is the number of points and $m$ is the number of dimensions.
A zonotope is a geometric shape formed by the Minkowski sum of line segments.
Definition 3 (Zonotope)
Given a center $c\in^{m}$ and generators $g_{1},\ldots,g_{p}$, a zonotope is
$\mathcal{Z}:=\left\{c+\sum_{i=1}^{p}\lambda_{i}g_{i}\;\middle|\;\lambda_{i}\in[-1,1],\forall i\in[1,\ldots,p]\right\}$
Zonotopes are computationally more efficient than convex hulls, with a construction complexity of $O(m\cdot p)$, where $m$ is the dimensionality and $p$ is the number of generators.
Finally, an interval is a simple shape defined by lower and upper bounds for each dimension, and it is equivalent to a multi-dimensional rectangle (or hyper-rectangle). Intervals are easy to construct with a complexity of $O(m)$, where $m$ is the dimensionality, and are often used in verification.
Definition 4 (Interval (aka Hyper-Rectangle))
Given a lower and upper bound $\underline{x},\overline{x}\in^{m}$ such that $\underline{x}_{(i)}\leq\overline{x}_{(i)}\forall i\in{1,\ldots,m}$, a multi-dimensional interval $\mathcal{I}\subset^{m}$ is
:=$\mathcal{I}:=\left\{x\in^{m}\;\middle|\;\underline{x}_{(i)}\leq x_{(i)}\leq\overline{x}_{(i)},\forall i\in[1,\ldots,m]\right\}$
Table 6 summarises the construction complexities of these different shapes. Ideally, convex hulls would be the preferred choice due to their precise and detailed representations of subspaces. However, their computational complexity renders them infeasible in high dimensions. Zonotopes provide a promising alternative, as they are more precise than hyper-rectangles while remaining computationally tractable. Despite their theoretical compatibility with complete verifiers, practical limitations arise because most state-of-the-art verifiers do not support zonotopes. Hyper-rectangles, or intervals, are the simplest to construct and are supported by all verifiers, making them the default choice in many verification pipelines.
Shape Construction Complexity (Big-O)
Convex Hull $O(p^{m/2})$
Zonotope $O(m\cdot p)$
Interval $O(m)$
Table 6: Construction complexity for different geometric shapes, where $m$ is the number of dimensions and $p$ is the number of points or generators.
3.4 Working with Embedding Spaces: Our Approach
We now formally define geometric and semantic subspaces of the embedding space. Our goal is to define subspaces on the embedding space m by using an effective algorithmic procedure. We will use notation $\mathcal{S}$ to refer to a subspace of the embedding space. Recall that an hyper-rectangle of dimension $m$ is a list of points $(a_{1},b_{1}),\ldots,(a_{m},b_{m})$ such that a point $x\in^{m}$ is a member if for every dimension $j$ we have $a_{j}\leq x_{j}\leq b_{j}$.